Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling 80 mph and train b is traveling 100 mph. Train A passes a station at 6:15 am. If train B passes the same station at 6:27 am, at what time will train B catch up to train A?
Found 2 solutions by mananth, mathmandotcom:
Answer by mananth(16949)


train B passes the station after 12 minutes
Train A speed is 80mph
so in 12 minutes Train A travels 12/60 *80
=16 miles.
..
train B speed = 100 mph. Train A speed speed 80 mph
so it catches up 20 miles per hour ( 100-80)
20 miles it catches up in 1 hour
16 miles it will catch up in ?
16/20 hours
4/5 hours.
4/5 * 60 minutes
=48 minutes.
..
Train B passes the station at 6.27 am
add 48 minutes to 6.27 am
it catches up at 7.15 am

Answer by mathmandotcom(31)

The solution to this problem lies in the equation d=rt
d=distance; da=distance train a travels, db=distance train b travels
r=rate; ra=80, rb=100
t=time; ta= time for train a, tb=time for train b
---------------------------------------------------------------------
da=db so ta*ra=tb*rb train A has a head start by 12 minutes or 80*12/60=16m
da=16+80t=100t=db
16+80t=100t subtract 80t
16=20t divide by 20
4/5=t in hours in minutes 4/5*60=48minutes; answer